# Beyond Joint Demonstrations: Personalized Expert Guidance for Efficient Multi-Agent Reinforcement Learning

### v1

@@ 4.2 Solving Personalized Expert-Guided MARL
Now, we can begin the development of a practical algorithm. By adopting Theorem 2 from Kang et al. (2018), we can substitute the JS divergence between occupancy measures with
@@
where $D(s_{i},a_{i})\mathrel{\mathop{\ordinarycolon}}\mathcal{S}_{i}\times\mathcal{A}_{i}\rightarrow(0,1)$ is a discriminative classifier to discern if the $(s_{i},a_{i})$ pair is from the demonstration or the current policy. However, this doesn’t account for the constraint in Equation equation 6. Therefore, by introducing an indicator function $\mathds{1}(s_{i},a_{i})$ for whether $(s_{i},a_{i})$ belongs to the domain of $\hat{\lambda}^{\pi_{\theta}}_{i}$ and $\hat{\lambda}^{\pi_{E_{i}}}$, we can obtain
@@
Since direct access to the transition distributions is unavailable, verifying whether each $(s_{i},a_{i})$ pair is within the domain is not directly feasible. To address this, we further approximate the indicator function using another discriminative classifier $D_{i}(s_{i},a_{i},s_{i}^{\prime})\mathrel{\mathop{\ordinarycolon}}\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\rightarrow(0,1)$, estimating the likelihood of a $(s_{i},a_{i},s_{i}^{\prime})$ tuple being from the demonstrations. In essence, it quantifies the likelihood that the corresponding $(s_{i},a_{i})$ induces a desired change in dynamics as observed in the demonstrations.
@@
6:         Update personalized dynamics discriminator $D_{\bar{\phi_{i}}}$:
@@
After approximating the inner JS divergence term in Equation equation 5 with two discriminators, we integrate the discriminative rewards derived from these discriminators into the environmental reward for the outer problem. Specifically, we estimate the reshaped reward $\hat{r}_{i}$ as $\hat{r}_{i}=r-\eta D_{\bar{\phi_{i}}}(s_{i},a_{i},s_{i}^{\prime})\log(1-D_{\phi_{i}}(s_{i},a_{i}))$, where $\phi_{i}$ and $\bar{\phi_{i}}$ are parameters for the two discriminators. The personalized behavior discriminator $D_{\phi_{i}}$ evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones, while the personalized dynamics discriminator $D_{\bar{\phi_{i}}}$ assesses if a local state-action pair induces a desired change in dynamics akin to that observed in the demonstration, adjusting the incentive weight accordingly. Subsequently, policy optimization is conducted by maximizing the long-term return with the reshaped reward. Our PegMARL algorithm, detailed in Algorithm 1, is compatible with any policy gradient methods, with MAPPO (Yu et al. (2021)) adopted in our implementation.
@@ 5 Experiments
5.1 Main Results with Personalized Demonstrations
We evaluate PegMARL with personalized demonstrations on both discrete gridworld environments (Figure 5) and a continuous multi-agent particle environment (Figure 7(a)). Our comparison includes four strong baselines: MAPPO (Yu et al. (2021)), a leading decentralized MARL algorithm; MAGAIL (Song et al. (2018)), a state-of-the-art multi-agent imitation learning algorithm; DM2 (Wang et al. (2023)), which combines distribution matching reward with environmental reward; and ATA (She et al. (2022)), one of the best multi-agent reward-shaping methods. Notably, MAGAIL and DM2 were not originally designed for personalized demonstrations. Hence, we have adapted them for use in personalized demonstration settings through necessary modifications.
(a) The lava scenario: the agents are homogeneous, aiming to reach corresponding diagonal positions without entering the lava.
@@ 5.1 Main Results with Personalized Demonstrations
(a) The lava scenario: the agents are homogeneous, aiming to reach corresponding diagonal positions without entering the lava.
@@
For each scenario, we collect two sets of personalized demonstrations for each agent: optimal and suboptimal. The average episodic rewards of suboptimal demonstrations are approximately half of their optimal counterparts (more details can be found in Appendix B). We execute each method in every environment with 10 distinct random initializations and plot the mean and variance across all runs. Oracle denotes the best possible return achievable with optimal policies.
@@
In contrast, PegMARL demonstrates superior generalizability across scenarios with more agents, maintaining stable performance despite suboptimal demonstrations. As DM2 has been adapted to accommodate personalized demonstrations, it closely resembles PegMARL but lacks the personalized dynamics discriminator component. As the number of agents increases, we would expect the number of ways for agents to interact to increase, and therefore for the personal demonstrations to be potentially misleading. Consequently, DM2 experiences a decrease in convergence speed as the number of agents increases. This highlights the importance of the personalized dynamics discriminator in PegMARL, enabling effective handling of inter-agent interactions in personalized demonstration scenarios, thereby ensuring its efficacy across varying agent counts.
How does PegMARL perform under the heterogeneous setting? The door scenario (Figure 5) contains two variants with varying levels of difficulty. The easy case is fairly straightforward: success should be achievable by adhering to personalized demonstrations, which illustrate how each agent navigates to their respective goal locations. As shown in Figure 7, most algorithms, except for MAGAIL, showcase proficient performance. Notably, PegMARL exhibits the swiftest convergence. We attribute MAGAIL’s failure in this case to its inability to direct the green agent to remain positioned at the green square — a behavior not explicitly demonstrated. This underscores the importance of environmental reward signals when integrating personalized demonstrations into multi-agent learning paradigms. The hard case necessitates a higher degree of agent cooperation: once the red agent gains entry to the middle room, the green agent must move aside from the green square to enable the red agent’s passage into the right room. In this complex setting, only PegMARL and DM2 demonstrate commendable convergence. PegMARL, equipped with the personalized dynamics discriminator, maintains faster convergence compared to DM2.
@@ 5.2 Results with Joint Demonstrations
While PegMARL is primarily designed to leverage personalized demonstrations, it can also be extended to utilize joint demonstrations. Unlike other MARL algorithms like DM2 (Wang et al. (2023)) that require compatible joint demonstrations sampled from co-trained policies to achieve convergence, PegMARL has the flexibility to utilize joint demonstrations even from non-co-trained policies. To demonstrate, we conduct comparative experiments within the StarCraft Multi-Agent Challenge (SMAC) environment (Samvelyan et al. (2019)). We employed the same two tasks that were used in DM2: 5mv6m, which involves 5 Marines (allies) against 6 Marines (enemies), and 3sv4z, which features 3 Stalkers (allies) versus 4 Zealots (enemies). The win rates of the demonstrations in both cases are approximately $30\%$.
@@
When the joint demonstrations are sampled from policies that are non-co-trained, the success rate of DM2 exhibits a significant decline in both tasks. In the 5mv6m task, DM2’s success rate drops to nearly 0, indicating a failure to learn. In contrast, PegMARL maintains a similar level of performance compared to results from co-trained demonstrations, with only a slight decrease in convergence speed. These findings underscore the versatility and effectiveness of PegMARL in scenarios requiring collaboration among diverse agents, demonstrating its robustness across various multi-agent environments.
@@ 6 Conclusions
In this work, we introduce PegMARL, a novel approach for Multi-Agent Reinforcement Learning, which adopts personalized expert demonstrations as guidance and allows agents to learn to cooperate. Specifically, the algorithm utilizes two discriminators to dynamically reshape the reward function: one provides incentives to encourage the alignment between policy behavior with provided demonstrations, and the other regulates incentives based on whether the behavior leads to the desired objective. We demonstrate PegMARL’s effectiveness, scalability, and robustness in both discrete and continuous environments, where it outperforms state-of-the-art decentralized MARL algorithms, pure imitation learning, and reward-shaping techniques. We observe that PegMARL can achieve near-optimal policies even with suboptimal demonstrations. Furthermore, We showcase PegMARL’s capability to leverage joint demonstrations and converge successfully, regardless of whether they are sampled from co-trained or non-co-trained policies.
@@ A.1 Discrete Gridworld Environments
We use two gridworld environments with discrete state and action space: the lava scenario and the door scenario (Figure 5). The agents in the lava scenario are homogeneous because they have the same objective: navigating to their corresponding goal location. The door scenario is heterogeneous: the assistant agent (green) must open the door while the other agent (red) must reach the goal.
@@ A.2 Cooperative Navigation
We modified the cooperative navigation task of multi-agent particle environment Lowe et al. (2017); Mordatch & Abbeel (2017) to evaluate the performance of our algorithm in a continuous environment. The modified environment consists of 2 agents, 2 goal landmarks and 1 wall between the agents and the goals. The agents need to navigate around the wall to occupy both landmarks. The state definition is the same as the original. We sparsify their original reward as follows:
@@
Appendix B Additional Experiments Details and Visualizations
Figure 10 shows an example of personalized demonstrations for the lava scenario. Table 1 summarizes the details of the suboptimal demonstrations, whose average episodic rewards are around 4.5 and about half of their optimal counterparts.
[Uncaptioned image] Figure 10: An example of personalized demonstrations for the lava scenario (we did not visualize all the optimal paths). There is only one agent in the environment. The agent may take both the left and the right path toward the goal.
Agent Id $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward 1 $\mathbb{R}^{2}$ 5 300 4.42 2 $\mathbb{R}^{2}$ 5 440 4.79 3 $\mathbb{R}^{2}$ 5 344 4.3 4 $\mathbb{R}^{2}$ 5 360 4.41 (a) The lava scenario Case Agent $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward easy red $\mathbb{R}^{2}$ 5 643 4.01 green $\mathbb{R}^{2}$ 5 612 4.39 hard red $\mathbb{R}^{2}$ 5 607 4.34 green $\mathbb{R}^{2}$ 5 593 4.36 (b) The door scenario
Table 1: The details of suboptimal demonstrations.

### v3

@@ 4.2 Solving Personalized Expert-Guided MARL
Now, we can begin the development of a practical algorithm. By adopting Theorem 2 from Kang et al. (2018) [^3] We refer the reader to Kang et al. (2018) for the full derivation. We use the same calculations with the change of notation $\rho_{\pi}\leftarrow\lambda_{i}^{\pi_{\theta}}$ and $\rho_{E}\leftarrow\lambda^{\pi_{E_{i}}}$., we can substitute the JS divergence between occupancy measures with
@@
where $D(s_{i},a_{i}):\mathcal{S}_{i}\times\mathcal{A}_{i}\rightarrow(0,1)$ is a discriminative classifier to discern if the $(s_{i},a_{i})$ pair is from the demonstration or the current policy. However, this doesn’t account for the constraint in equation 6. Therefore, by introducing an indicator function $\mathds{1}(s_{i},a_{i})$ for whether $(s_{i},a_{i})$ belongs to the domain of $\hat{\lambda}^{\pi_{\theta}}_{i}$ and $\hat{\lambda}^{\pi_{E_{i}}}$, we can obtain
@@
Since direct access to the transition distributions is unavailable, verifying whether each $(s_{i},a_{i})$ pair is within the domain is not directly feasible. To address this, we further approximate the indicator function using another discriminative classifier $D_{i}(s_{i},a_{i},s_{i}^{\prime}):\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\rightarrow(0,1)$, estimating the likelihood of a $(s_{i},a_{i},s_{i}^{\prime})$ tuple being from the demonstrations. In essence, it quantifies the likelihood that the corresponding $(s_{i},a_{i})$ induces a desired change in local state transition as observed in the demonstrations, effectively serving as a learned proxy for the $\epsilon$-constrained domain in Equation 6.
@@
6:         Update personalized transition discriminator $D_{\bar{\phi_{i}}}$:
@@
After approximating the inner JS divergence term in equation 5 with two discriminators, we integrate the discriminative rewards derived from these discriminators into the environmental reward for the outer problem. Specifically, we estimate the reshaped reward $\hat{r}_{i}$ as $\hat{r}_{i}=r-\eta D_{\bar{\phi_{i}}}(s_{i},a_{i},s_{i}^{\prime})\log(1-D_{\phi_{i}}(s_{i},a_{i}))$, where $\phi_{i}$ and $\bar{\phi_{i}}$ are parameters for the two discriminators. The personalized behavior discriminator $D_{\phi_{i}}$ evaluates local state-action pairs, providing positive incentives for actions that align with the demonstration and negative incentives for divergent ones, while the personalized transition discriminator $D_{\bar{\phi_{i}}}$ assesses if a local state-action pair induces a desired transition in local state akin to that observed in the demonstration, adjusting the incentive weight accordingly. Subsequently, policy optimization is conducted by maximizing the long-term return with the reshaped reward.
Our PegMARL algorithm, detailed in Algorithm 1, is compatible with any policy gradient methods, with MAPPO (Yu et al. (2021)) adopted in our implementation. While presented in a fully observable setting, PegMARL can be adapted for partially observable scenarios. In such cases, we process observations from joint environment rollouts by removing dimensions not observable in the Personalized MDP for discriminators inputs. Notably, PegMARL can also utilize joint demonstrations; in this case, the observation processing step is unnecessary as both demonstrations and environment rollouts are derived from the same joint MDP. Compared to DM2 (Wang et al., 2023), which also uses a discriminator-based approach for multi-agent learning from demonstrations, PegMARL introduces the personalized transition discriminator $D_{\bar{\phi_{i}}}$ as a key innovation. This additional discriminator helps address the challenge illustrated in Figure 3 by evaluating whether local state-action pairs induce desired transitions similar to those in the demonstrations, effectively filtering out less desirable actions while retaining beneficial ones. This component enables PegMARL to handle both personalized demonstrations and joint demonstrations from various sources, including those sampled from non-co-trained policies, while DM2 requires demonstrations from co-trained policies to achieve convergence.
@@ 5 Experiments
(a) The lava scenario: the agents are homogeneous, aiming to reach corresponding diagonal positions without entering the lava. The episode ends if any agent steps into the lava.
@@
5.1 Main Results with Personalized Demonstrations
We evaluate PegMARL with personalized demonstrations on discrete gridworld environments (Figure 5) and a continuous multi-agent particle environment (MPE, Figure 8(a)). The gridworld environment is fully observable with discrete state and action spaces, while the MPE is partially observable with a continuous state space and discrete action space. Both environments feature sparse reward signals, which adds to the challenge for standard MARL algorithms to learn effectively.
Our comparison includes four strong baselines: MAPPO (Yu et al. (2021)), a leading decentralized MARL algorithm; MAGAIL (Song et al. (2018)), a state-of-the-art multi-agent imitation learning algorithm; DM2 (Wang et al. (2023)), which combines distribution matching reward with environmental reward; and ATA (She et al. (2022)), one of the best multi-agent reward-shaping methods. Notably, MAGAIL and DM2 were not originally designed for personalized demonstrations. Hence, we have adapted them for use in personalized demonstration settings through necessary modifications. Specifically, we processed the observations from joint environment rollouts by removing dimensions not observable in the Personalized MDP, similar to our approach in PegMARL. This adjustment ensures that the inputs to their respective discriminators are compatible with the personalized demonstration data. In this context, MAGAIL serves as an ablation of PegMARL lacking the environmental signal and transition discriminator, while DM2 represents an ablation lacking only the transition discriminator. Details on algorithm implementation and hyperparameter choices can be found in Appendix C.
@@ 5.1 Main Results with Personalized Demonstrations
For each scenario, we collect two sets of personalized demonstrations for each agent: optimal and suboptimal, by training single-agent RL in simplified environments that focus on each agent’s specific tasks. The average episodic rewards of suboptimal demonstrations are approximately half of their optimal counterparts (more details can be found in Appendix B). We execute each method in every environment with 10 distinct random initializations and plot the mean and variance across all runs. Oracle denotes the best possible return achievable with optimal policies.
@@
In contrast, PegMARL demonstrates superior generalizability across scenarios with more agents, maintaining stable performance despite suboptimal demonstrations. As DM2 has been adapted to accommodate personalized demonstrations, it closely resembles PegMARL but lacks the personalized transition discriminator component. As the number of agents increases, we would expect the space of possible interactions among agents to expand, and therefore for the personal demonstrations to be potentially misleading. Consequently, DM2 experiences a decrease in convergence speed as the number of agents increases. This highlights the importance of the personalized transition discriminator in PegMARL, enabling effective handling of inter-agent interactions in personalized demonstration scenarios, thereby ensuring its efficacy across varying agent counts.
How does PegMARL perform under the heterogeneous setting? The door scenario (Figure 5) contains two variants with varying levels of difficulty. The easy case is fairly straightforward: success should be achievable by adhering to personalized demonstrations, which illustrate how each agent navigates to their respective goal locations. As shown in Figure 7, most algorithms, except for MAGAIL, showcase proficient performance. Notably, PegMARL exhibits the swiftest convergence. We attribute MAGAIL’s failure in this case to its inability to direct the green agent to remain positioned at the green square — a behavior not explicitly demonstrated. This underscores the importance of environmental reward signals when integrating personalized demonstrations into multi-agent learning paradigms. The hard case necessitates a higher degree of agent cooperation: once the red agent gains entry to the middle room, the green agent must move aside from the green square to enable the red agent’s passage into the right room. In this complex setting, only PegMARL and DM2 demonstrate commendable convergence. PegMARL, equipped with the personalized transition discriminator, maintains faster convergence compared to DM2.
@@ 5.2 Results with Joint Demonstrations
While PegMARL is primarily designed to leverage personalized demonstrations, it can also be extended to utilize joint demonstrations. Unlike other MARL algorithms like DM2 (Wang et al. (2023)) that require compatible joint demonstrations sampled from co-trained policies to achieve convergence, PegMARL has the flexibility to utilize joint demonstrations even from non-co-trained policies, which potentially contain conflicting behaviors. To demonstrate, we conduct comparative experiments within the StarCraft Multi-Agent Challenge (SMAC) environment (Samvelyan et al. (2019)). We adopted DM2’s original implementation and parameters for accurate results replication. We employed the same two tasks that were used in DM2: 5mv6m, which involves 5 Marines (allies) against 6 Marines (enemies), and 3sv4z, which features 3 Stalkers (allies) versus 4 Zealots (enemies). The win rates of the demonstrations in both cases are approximately $30\%$.
@@
When the joint demonstrations are sampled from policies that are non-co-trained, the success rate of DM2 exhibits a significant decline in both tasks. In the 5mv6m task, DM2’s success rate drops to nearly 0, indicating a failure to learn. In contrast, PegMARL maintains a similar level of performance compared to results from co-trained demonstrations, with only a slight decrease in convergence speed. These findings underscore the versatility and effectiveness of PegMARL in scenarios requiring collaboration among diverse agents, demonstrating its applicability and robustness across various multi-agent scenarios and demonstration types.
@@ 6 Conclusions
In this work, we introduce PegMARL, a novel approach for Multi-Agent Reinforcement Learning, which adopts personalized expert demonstrations as guidance and allows agents to learn to cooperate. Specifically, the algorithm utilizes two discriminators to dynamically reshape the reward function: one provides incentives to encourage the alignment between the policy behavior with provided demonstrations, and the other regulates incentives based on whether the behavior leads to the desired objective. We demonstrate PegMARL’s effectiveness, scalability, and robustness in both discrete and continuous environments, where it outperforms state-of-the-art decentralized MARL algorithms, pure imitation learning, and reward-shaping techniques. We observe that PegMARL can achieve near-optimal policies even with suboptimal demonstrations. Furthermore, We showcase PegMARL’s capability to leverage joint demonstrations and converge successfully, regardless of whether they are sampled from co-trained or non-co-trained policies.
Applicability and Limitations: PegMARL is most applicable when: a) The task allows for some degree of independent decision-making, allowing individual agent behaviors to be isolated and demonstrated, even if only partially. b) The environment provides feedback on cooperative behaviors, enabling agents to learn coordination beyond individual demonstrations. For example, in the door scenario, agents gain exploration guidance for door operation and navigation from personalized demonstrations. At the same time, environmental feedback on joint success drives them to learn coordination about timing and positioning (e.g., the green agent learning to make way for the red agent). Similarly, in the lava scenario, personalized demonstrations provide guidance for lava-avoiding paths. Still, environmental penalties for collisions and rewards for collective goal achievement encourage the emergence of cooperative navigation strategies.
However, PegMARL may face challenges in tasks requiring continuous, seamless collaboration among agents, such as cooperative object lifting and relocation by multiple agents. In these scenarios, where constant, integrated teamwork is crucial, the reliance on personalized demonstrations may not provide sufficient guidance for complex, ongoing coordination. Despite this, PegMARL’s capability to utilize joint demonstrations from diverse sources, including non-co-trained policies, suggests the potential for addressing this limitation. It would be interesting future work to explore PegMARL’s capabilities to handle more intricate and continuously collaborative tasks.
Another limitation of our work is the absence of formal convergence guarantees. While DM2 establishes convergence for individual agents optimizing distribution matching objectives, our approach introduces additional complexity through the transition discriminator’s dynamic adjustment of demonstration influence. The interplay between the dual discriminators and multi-agent dynamics poses challenges for theoretical analysis, particularly since the joint distribution cannot be directly recovered from individual agent demonstrations in our setting. While our extensive empirical results demonstrate robust convergence across diverse scenarios, establishing theoretical convergence guarantees remains an interesting direction for future work.
@@ A.1 Discrete Gridworld Environments
We use two gridworld environments with discrete state and action space: the lava scenario and the door scenario (Figure 5). The agents in the lava scenario are homogeneous because they have the same objective: navigating to their corresponding goal location. The door scenario has heterogeneous agents: the assistant agent (green) must open the door while the other agent (red) must reach the goal.
@@ A.2 Cooperative Navigation
We modified the cooperative navigation task of multi-agent particle environment Lowe et al. (2017); Mordatch & Abbeel (2017) to evaluate the performance of our algorithm in a continuous environment. The modified environment consists of 2 agents, 2 goal landmarks and 1 wall between the agents and the goals. The agents need to navigate around the wall to occupy both landmarks. Each agent has a partial observation of the environment, including its position and velocity as well as the relative positions of other agents and landmarks. The action space is discrete, consisting of moving left, right, up, down, and staying in place. We sparsify their original reward as follows:
@@
A.3 The SMAC Environment
The StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. (2019)) is an environment specifically designed for evaluating multi-agent reinforcement learning (MARL) algorithms. Built on the StarCraft II game, SMAC presents a variety of challenging scenarios where multiple agents must collaborate to control individual units in combat. Each agent has partial observation of its surroundings within a certain sight range. The action space is discrete, including move[direction], attack[enemy_id], stop, and no-op. A global reward is given at each timestep based on the total damage dealt to the enemy units, with an additional bonus of 10 or 200 awarded for killing each enemy unit or for winning a combat respectively. We consider the same two tasks as used in DM2 (Wang et al. (2023)):
  • 5m_vs_6m: The allied team consists of 5 Marines and the enemy team controls 6 Marines.
  • 3s_vs_4z: The allied team controls 3 stalker units and faces 4 enemy Zealot units.
Appendix B Demonstration Details
B.1 Personalized Demonstrations
We let each individual agent perform their designated task in the same environment map without the other agents’ presence to collect personalized demonstrations. This can be implemented through two main methods: a) recording expert human demonstrations for each agent type, and b) training single-agent RL in simplified environments that focus on their specific tasks. In our experiments, we adopted the latter option. It takes about 2 hours to train a single-agent PPO for each personalized task.
Discrete Gridworld Environments: Figure 10 shows an example of personalized demonstrations for the lava scenario, and Figure 2 visualizes the personalized demonstrations for the door scenario. We summarize the details of the suboptimal demonstrations for the gridworld environment in Table 1, where the average episodic rewards are approximately 4.5, about half of their optimal counterparts.
Figure 10: An example of personalized demonstrations for the lava scenario (we did not visualize all the optimal paths). There is only one agent in the environment. The agent may take both the left and the right path toward the goal.
Agent Id $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward
1 $\mathbb{R}^{2}$ 5 300 4.42
2 $\mathbb{R}^{2}$ 5 440 4.79
3 $\mathbb{R}^{2}$ 5 344 4.3
4 $\mathbb{R}^{2}$ 5 360 4.41
(a) The lava scenario
Case Agent $\mathcal{S}$ $\mathcal{A}$ Samples Average Episodic Reward
easy red $\mathbb{R}^{2}$ 5 643 4.01
green $\mathbb{R}^{2}$ 5 612 4.39
hard red $\mathbb{R}^{2}$ 5 607 4.34
green $\mathbb{R}^{2}$ 5 593 4.36
(b) The door scenario
Table 1: The details of suboptimal demonstrations for the gridworld environment.